OpenStudy (anonymous):

jerome walked 2/3 miles in the morning and 11/12 miles in the evening how many miles did he walk

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }+\frac{ 11 }{ 12 }=?\]

OpenStudy (anonymous):

Use LCD. \[\frac{ 8 }{ 12 }+\frac{ 11 }{ 12 }=?\]